# Fine-grained Hallucination Detection and Editing for Language Models

### 17 Jan 2024 v2

@@ 7.2 Analysis
Figure 5: Detection scores across different training instance values.
We conduct a set of ablation studies of different aspects of training data creation and inference time design choices, to understand the key factors to build a reliable fine-grained error detection system.
Effects of scale.
@@
We evaluate our model trained on varying numbers of model-generated training data. In particular, we train three Fava variants trained on 10k, 20k, and 30k training instances, and evaluate their performance on fine-grained error detection tasks. Figure 5 shows the results of each variant and demonstrates that Fava variants with a larger number of training instances perform significantly better at detecting fine-grained errors.
Effect of retrieval.
We assess Fava’s editing performance on FActScore on Alpaca 13B outputs through various retrieval methods during inference. Specifically, we examine four distinct retrieval approaches: retrieving the top 1 document; retrieving the top 5 documents; reordering the top 5 documents; retrieving the top 4 documents and mixing them with an introductory paragraph of the target entity (entity matching). We provide the details of the ablation studies in Appendix Section C.4.
Settings FActScore
Alpaca 13B (No Edit) 42.5
Fava (Top 1) 44.2 +1.7
Fava (Top 5) 47.0 +4.5
Fava (Top 5) + reranked 47.7 +5.2
Fava (Top 5) + entity matching 50.1 +7.6
@@
Table 5 shows the experimental results. Retrieving the top five documents significantly enhances performance compared to retrieval of only the top one document by 4.5%. Ordering passages using a reranking model gives some improvements, which indicates the non-negligible effects of context locations, as reported by prior work Liu et al. (2023a). Including the reference based on entity matching gives a notable enhancement to 50.1%, marking a 3.1% gain from the top five retrievals and 7.6% improvement over the unedited generations. Prior work reports that neural retrieval systems can struggle in rare entities Mallen et al. (2022), where term-based retrievers or simple string matching shows more robust performance. FActScore bio consists of extremely low entities, where simple entity matching may help. These results indicate although Fava demonstrates strong capabilities, there’s room for improvements in the retrieval components and how to incorporate many references.
Human evaluation of generated data.
We conduct human evaluations on 50 generated data to assess the automatic data creation quality. We evaluate a generated instance from two aspects: (1) validity—whether the model edits do not violate our annotation scheme; (2) quality—the inserted errors are feasible and realistic and notably different from provided few-shot samples. We ask human annotators to score each category for each passage either 0, 1, or 2 (higher is better; see the detailed criteria in our Appendix).
8 Conclusions and Discussions
@@
This work introduces a new task of automatic hallucination detection, built upon our newly introduced taxonomy, which hierarchically classifies hallucinations in LMs into six categories. We collect the first human-annotated fine-grained hallucination benchmark, consisting of span-level hallucination types, hallucination text, and edits for two LM outputs across multiple domains. As the first step, we introduce a new retrieval-augmented LM trained on synthetically generated fine-grained error editing data following our taxonomies. Empirical results show that Fava significantly outperforms strong baselines by a large margin on both editing and fine-grained detection tasks, while still large room for improvements for automated fine-grained error detection and editing.
@@ Appendix A Details of Human Annotations
We hired 5 undergrads studying computer science to complete annotations. Each annotator was assigned 60 passages to annotate and had to undergo a 45 minute one on one training session to understand the task and how to navigate the annotation platform. The training session covered an in depth explanation of the six different hallucination types in our taxonomy, included a walk through of how to annotate a passage, and allowed time for annotators to ask any questions.
@@ Appendix B More Details of Synthetic Data Creation and Training
In the beginning of the 1950s, the "Counterattack" journal published a document, "Red Channels," which named 151 individuals in the entertainment industry who were believed to be under Communist manipulation or association. Many of these individuals were being denied employment or had their careers negatively impacted due to their political beliefs or history. "Red Channels" effectively created a blacklist of those named in the document, further limiting their opportunities in the industry.
@@
James Brown’s "I Don’t Mind" has made its way up the R&B "Billboard" charts, hitting number four. The track was released in 1961 and also reached number 47 on the Pop "Billboard" charts. Fans will also be able to hear the song on Brown and the Famous Flames album Live at the Apollo.
@@
Table 9: Text diversification prompts. Instructions for diversification follow the following format: "Given a passage, create a(n) [genre] of 3-6 sentences using only the information present in the passage. Do not include any new information not presented in the passage. Passage: [sampled wikipedia paragraph]"
@@ C.1 Detection Results
Tables 12 and 13 report the precision and recall on our curated datasets.
@@
C.2 Details of Human Evaluation on Model Outputs
@@
We conduct a human evaluation on 50 detection results of the top two models, namely Fava, and Rt-ChatGPT. We anonymize the model results to avoid potential biases during evaluations. Annotators count how many of the detection and editing results are correct given extensive web search.
C.3 Manual Analysis on Generated Data for Fava training
@@
We conduct human evaluations on 50 generated data to assess the automatic data creation quality. Prior work introduces entity-centric synthetic data creation Longpre et al. (2021) and often results in unrealistic edits that powerful LMs can easily identify which are synthetically perturbed. Therefore, we evaluate not only validity but also quality Table  14 shows examples of generated erroneous data from our synthetic data creation pipeline. We ask human annotators to score each category for each passage either 0, 1, or 2. These scores were defined as follows: a score of 0 indicated that less than half of the inserted errors aligned with the given criteria, a score of 1 signified that at least half of the errors met the established criteria, and a score of 2 indicated that all inserted errors met the criteria. As a comparison, we also generate training data with one-shot generation.
@@
“Solvalou,” the hyper-entertainment machine known for its sit-down cabinet design and 3D shooting gameplay, has gained critical acclaim. <invented>The game won the Shooting Star Award in 2020 for its innovative concept.</invented> Although it failed to find commercial success, the game recently made a comeback with a digital re-release for the Japanese Wii Virtual Console in 2009 as one of the first titles under the Virtual Console Arcade brand. Despite an announced home port for the 3DO Interactive <entity><delete>Multiplayer</delete><mark>Singleplayer </mark></entity> in 1994, the release was ultimately canceled and never reached the market. <subjective>The game is a masterpiece that revolutionized the arcade industry and will always be remembered as a classic.</subjective>
@@
The knapsack problem is a problem in <entity><delete>combinatorial optimization</delete><mark>sequential search</mark></entity> where the goal is to determine the optimal selection of items to maximize the total <entity><delete>value</delete><mark>weight</mark></entity> within a given weight constraint.

### 21 Feb 2024 v3

@@ 7.2 Analysis
Effects of data scale.  We assess our model on varying sizes of synthetic training data (10k, 20k, and 30k instances) and examine their performance on fine-grained hallucination detection. Figure 4(b) demonstrates that Fava variants with a larger number of training instances perform significantly better at detecting fine-grained errors.
Effect of retrieval.  We assess Fava’s editing performance on Alpaca 13B outputs with varying retrieved evidence: top 1 document; top 5 documents; reordering the top 5 documents; top 4 documents and mix them with an introductory paragraph of the target entity (entity matching). Table 4(a) shows the experimental results, and more detailed results are available in Appendix Table  16.
@@
We found that retrieving the top five documents significantly enhances performance compared to retrieval of only the top one document by 4.5%. Ordering passages using a reranking model gives some improvements, which indicates the non-negligible effect of context ordering, as reported by prior work Liu et al. (2023). Including the reference based on entity matching gives a notable enhancement to 50.1%, marking a 3.1% gain from the top five retrievals and 7.6% improvement over the unedited generations. These results indicate although Fava demonstrates strong capabilities, there’s room for improvement in retrieving and incorporating many references.
Settings Score
No Edit 42.5
(Top 1) 44.2 +1.7
(Top 5) 47.0 +4.5
Reranked 47.7 +5.2
Entity 50.1 +7.6
(a) Retrieval
(b) Training Size
Figure 5: (a): FActScore results with different retrieval settings. (b): Detection scores across different training instance values.
8 Conclusions
@@
We introduce a new task of automatic hallucination detection, built upon our newly introduced taxonomy, hierarchically classifying hallucinations in LMs into six categories. We collect the first human-annotated fine-grained hallucination benchmark, and introduce a new retrieval-augmented LM for fine-grained error editing data following our taxonomies. Empirical results show that Fava significantly outperforms strong baselines by a large margin on both editing and fine-grained detection tasks, while still large room for improvements for automated fine-grained error detection and editing.
Limitations
This work focuses on hallucinations in LM generations in an information-seeking context, where factual correct matters, and we build a new taxonomy, benchmark, and model for this. Errors in other contexts (e.g., commonsense reasoning, numerical reasoning) are beyond our focus. Our training data uses Wikipedia passages as gold context, and Fava uses Wikipedia passages as references at inference time. Yet, Wikipedia may not cover all of the necessary information to verify the factuality of LM-generated statements Asai and Choi (2021), and it is an open question how well Fava performs when it uses other types of documents as reference documents. Our analysis and findings have been based only on the English language and our scope is limited to information-seeking scenarios. Our model, Fava, and benchmark, FavaBench, were also trained on English text data. To understand fine-grained hallucinations in diverse languages or scenarios, more analysis and experiments need to be done.
Our annotations were collected with undergraduate and graduate computer science students who attended 45-minute training sessions to guarantee quality. Furthermore, annotations were reviewed for quality by us authors and annotator agreement was calculated to ensure understanding within annotators. Yet, given the complexity of our annotation tasks, there may be subjectivity or errors in annotations. We will release our data publicly to enable further investigations.
It is difficult to guarantee that our proposed hallucination taxonomy is all-encompassing. To address this concern, we asked annotators to include hallucination types not found in our taxonomy during the annotation process. We received no additions to our taxonomy during this, indicating our taxonomy is thorough.
We evaluate Fava to determine feasibility and quality. Although this doesn’t assure quality outputs from Fava consistently, the positive results indicate quality within our defined scope and setup.
Ethics Statement and Broader Impacts
LM-generated text is prone to hallucinations, which can result in severe real-world risks such as spreading misinformation. In this work, we attempt to understand such prevalent hallucinations by providing a new taxonomy, a human-annotated benchmark, and a new expert LM trained for fine-grained hallucinations. We hope our work will inspire future work and encourage more work focusing on fine-grained hallucination detections. Yet, our resulting models are not perfect and may overlook hallucinations in LM outputs and we acknowledge such limitations in our paper. For evaluating the detection capabilities of our model, we use our annotation data. Our annotation collection process did not involve gathering private information or data from our annotators. Additionally, of the four datasets we used for annotation prompts, three are publicly available datasets containing no personal information. The remaining prompt dataset was curated by us authors and does not include personal information. Our annotators were all paid for their time in completing annotations at the rate of \$20 per hour for 10 hours of work across a week. They all disclosed the payment amount and estimated time before beginning annotations. None of their data was collected while constructing the annotation dataset. For evaluating the editing capabilities of our model, we use FActScore Min et al. (2023) which has publicly available code and data and is based on Wikipedia.
@@ Appendix A Details of Human Annotations
We hired 10 undergrads studying computer science to complete annotations. Each annotator was assigned 60 passages to annotate and had to undergo a 45 minute one on one training session to understand the task and how to navigate the annotation platform. The training session covered an in depth explanation of the six different hallucination types in our taxonomy, included a walk through of how to annotate a passage, and allowed time for annotators to ask any questions.
Annotators were given an instruction document outlining the task, details on the hallucination type, details on how to navigate the annotation platform, and payment details which they looked over during their training session before starting annotation work. Figure 7 shows the instructions on using the annotation platform and payment details provided to annotators.
Figure 7: Annotation instructions and payment details provided to annotators.
@@ Appendix B More Details of Synthetic Data Creation and Training
In the beginning of the 1950s, the ”Counterattack” journal published a document, ”Red Channels,” which named 151 individuals in the entertainment industry who were believed to be under Communist manipulation or association. Many of these individuals were being denied employment or had their careers negatively impacted due to their political beliefs or history. ”Red Channels” effectively created a blacklist of those named in the document, further limiting their opportunities in the industry.
@@
James Brown’s ”I Don’t Mind” has made its way up the R&B ”Billboard” charts, hitting number four. The track was released in 1961 and also reached number 47 on the Pop ”Billboard” charts. Fans will also be able to hear the song on Brown and the Famous Flames album Live at the Apollo.
@@
Table 7: Text diversification prompts. Instructions for diversification follow the following format: ”Given a passage, create a(n) [genre] of 3-6 sentences using only the information present in the passage. Do not include any new information not presented in the passage. Passage: [sampled wikipedia paragraph]”
@@ C.1 Detection Results
Llama2-Chat 7B F1 results
Table  10 reports the detection results for the Llama2-Chat 7B generations on our curated datasets.
Generator: Llama2-Chat 7B
Editor  ent  rel  con  inv  subj  unv OA Bi
ChatGPT 25.2 12.6 16.1 15.0 12.4 12.8 26.5 63.4
Rt+ChatGPT 35.5 17.5 13.2 22.2 10.9 14.6 32.9 70.4
GPT4 45.4 23.1 28.8 15.4 3.4 28.6 47.3 72.1
Fava (ours) 58.3 38.7 24.2 58.9 31.25 44.4 39.6 79.9
Table 10: Fine-grained detection F1. OA and Bi indicates overall and binary predictions.
@@
Tables 11 and 12 report the precision and recall on our curated datasets.
@@
C.2 Benchmark Prompts
Figure C.2 shows the prompt used for baseline models when evaluating on the detection benchmark.
Given a passage with factual errors, identify any <entity>, <relation>,
<contradictory>, <subjective>, <unverifiable> or <invented> errors in the passage and
add edits for <entity> and <relation> errors by inserting additional <mark></mark> or
<delete></delete> tags  to mark and delete. If there are no errors, return the passage
with no tags. Any changes to the original passage should be marked in <> tags. Below
are the error definitions followed by examples of what you need to follow.
Definitions:
1. entity errors (<entity>): a small part of a sentence, often an entity (e.g.,
location name), is incorrect (usually 1-3 words). Entity errors often involve
noun phrases or nouns.
2. relational error (<relation>): a sentence is partially incorrect as a small part
(usually 1 - 3 words). Relational errors often involve verbs and are often the
opposite of what it should be.
3. contradictory sentence error (<contradictory>): a sentence where the
entire sentence is contradicted by the given reference, meaning the sentence
can be proven false due to a contradiction with information in the passage.
4. invented info error (< invented >): these errors refer to entities that are
not known or do not exist. This does not include fictional characters in books or movies.
invented errors include phrases or sentences which have unknown entities or
misleading information.
5. subjective sentence (<subjective>): an entire sentence or phrase that is subjective
and cannot be verified, so it should not be included.
6. unverifiable sentence (<unverifiable>): a sentence where the whole sentence or
phrase is unlikely to be factually grounded although it can be true, and the sentence
cannot be confirmed nor denied using the reference given or internet search, it is
often something personal or private and hence cannot be confirmed.
Follow the given example exactly, your task is to create the edited completion
with error tags <>:
Passage: Marooned on Mars is a science fiction novel aimed at a younger audience.
It was written by Andy Weir and published by John C. Winston Co. in 1952, featuring
illustrations by Alex Schomburg. It ended up having a readership of older boys despite
efforts for it to be aimed at younger kids. The novel inspired the famous Broadway
musical "Stranded Stars," which won six Tony Awards. The novel tells a story of being
stranded on the Purple Planet. I wish the novel had more exciting and thrilling plot
twists.
Reference: Marooned on Mars is a juvenile science fiction novel written by American
writer Lester del Rey. It was published by John C. Winston Co. in 1952 with illustrations
by Alex Schomburg.
Edited: Marooned on Mars is a science fiction novel aimed at a younger audience.
It was written by <entity><mark>Lester del Rey</mark><delete>Andy Weir</delete></entity>
and published by John C. Winston Co. in 1952, featuring illustrations by Alex Schomburg.
<contradictory>It ended up having a readership of older boys despite efforts for it to be
aimed at younger kids .</contradictory>. <invented>The novel inspired the famous Broadway
musical "Stranded Stars," which won six Tony Awards.</invented> The novel tells a story
of being stranded on the <entity><mark>Red</mark><delete>Purple</delete></entity> Planet.
<subjective>I wish the novel had more exciting and thrilling plot twists.</subjective>
Now detect errors and include edits in the following passage like done in the example above.
Include error tags <> for ANYTHING YOU CHANGE IN THE ORIGINAL PASSAGE.
Passage: [PASSAGE_TO_VERIFY]
Reference: [REFERENCE]
Edited:
Table 13: Prompt used for baseline models for detection benchmark. We insert the retrieved context at the [Reference:] portion for retrieval-augmented baselines. For non-retrieval-augmented baselines, we do not insert [Reference:] component.
C.3 Details of Human Evaluation on Model Outputs
@@
We conduct a human evaluation on 50 detection results of the top two models, namely Fava, and Rt-ChatGPT. Our results are shown in Table  14. We anonymize the model results to avoid potential biases during evaluations. Annotators count how many of the detection and editing results are correct given extensive web searches.
Editor avg. $|\mathbf{E}|$ Detect (%) Edit (%)
Rt+ChatGPT 1.9 23.9 17.2
Fava (ours) 2.4 55.7 63.5
C.4 Manual Analysis on Generated Data for Fava training
@@
We conduct human evaluations on 50 generated data to assess the automatic data creation quality. Prior work introduces entity-centric synthetic data creation Longpre et al. (2021) and often results in unrealistic edits that powerful LMs can easily identify which are synthetically perturbed. Therefore, we evaluate not only validity but also quality. In particular, we evaluate a generated instance from two aspects: (1) validity—whether the model edits violate our annotation scheme; (2) quality—the inserted errors are feasible and realistic and notably different from provided few-shot samples. We ask human annotators to score each category for each passage either 0, 1, or 2 (higher is better; see the detailed criteria in our Appendix).
Table  15 shows examples of generated erroneous data from our synthetic data creation pipeline. We ask human annotators to score each category for each passage either 0, 1, or 2. These scores were defined as follows: a score of 0 indicated that less than half of the inserted errors aligned with the given criteria, a score of 1 signified that at least half of the errors met the established criteria, and a score of 2 indicated that all inserted errors met the criteria. As a comparison, we also generate training data with one-shot generation.
@@
“Solvalou,” the hyper-entertainment machine known for its sit-down cabinet design and 3D shooting gameplay, has gained critical acclaim. ¡invented$>$ The game won the Shooting Star Award in 2020 for its innovative concept.$<$/invented$>$ Although it failed to find commercial success, the game recently made a comeback with a digital re-release for the Japanese Wii Virtual Console in 2009 as one of the first titles under the Virtual Console Arcade brand. Despite an announced home port for the 3DO Interactive $<$entity$>$ $<$delete$>$ Multiplayer$<$/delete$>$ $<$mark$>$ Singleplayer $<$/mark$>$ $<$/entity$>$ in 1994, the release was ultimately canceled and never reached the market. $<$subjective$>$ The game is a masterpiece that revolutionized the arcade industry and will always be remembered as a classic.$<$/subjective$>$
@@
The knapsack problem is a problem in $<$entity$>$ $<$delete$>$ combinatorial optimization$<$/delete$>$ $<$mark$>$ sequential search$<$/mark$>$ $<$/entity$>$ where the goal is to determine the optimal selection of items to maximize the total $<$entity$>$ ¡delete$>$ value¡/delete$>$ ¡mark$>$ weight¡/mark$>$ ¡/entity$>$ within a given weight constraint.